OpenStudy (anonymous):

Greg Lawrence anticipates he will need approximately $218,700 in 10 years to cover his 3-year-old daughter’s college bills for a 4-year degree. How much would he have to invest today, at an interest rate of 7 percent compounded semiannually? (Please use only the following provided Table) (Round your answer to 2 decimal places. Omit the "$" sign in your response.)

OpenStudy (anonymous):

Er, can you use a calculator?

OpenStudy (anonymous):

Well, if you can, then here's the answer : \[2,18,700 = P ( 1+ 7/200)^{8}\] Where P= Principal / The amount Greg has to invest today. x = 218700 / 1.316 x = 166185.4 ( Approx)